A triangle has an area of 72 square inches. If the base of the triangle has a length of 18 inches, what is the height of the triangle? Use the formula for the area of a triangle:
Area = 1/2(base)(height)

Answer by macston(5194)

Area=(1/2)(base)(height) Multiply each side by 2.
(2)(Area)=(base)(height) Divide each side by base.
(2)(Area)/(base)=height
(2)(72sqin)/(18in)=height
(2)(4in)=height
8in=height ANSWER: The height is 8 inches.
CHECK:
Area=1/2(base)(height)
72 sq in=1/2(18 in)(8 in)
72 sq in=(9 in)(8 in)
72 sq in=72 sq in